OpenStudy (anonymous):

What is the quotient (5-x)/(x^2+3x-4) divided by (x^2-2x-15)/(x^2+5x+4) in simplified form? State any restrictions on the variable

OpenStudy (anonymous):

\[\frac{ 5-x }{ x^{2} +3x-4}\div \frac{ x^{2}-2x-15 }{ x^{2}+5x+4 }\]

OpenStudy (anonymous):

Factorise the quadratic expressions. Change the division sign to multiplication by taking reciprocal of the 2nd part of equation. See if things cancel out

OpenStudy (anonymous):

Uh, what? Could you pretty please break it down really simple?

OpenStudy (trojanpoem):

-(x-5)/ (x + 4) ( x - 1) * (x + 4) ( x + 1)/ (x - 5) ( x + 3) (x+4) out and (x - 5) out -(x+1)/(x+3)(x-1) Mathes on pc = hell

OpenStudy (anonymous):

okay. Can you factorise all the quadratic expressions?

OpenStudy (anonymous):

@TrojanPoem tell him/her how to do it. Not the answer...

OpenStudy (trojanpoem):

I thought , if he saw it answered , he would understand how it's done.

OpenStudy (anonymous):

*she And I would like to have the process rather than just the answer, thank you though.

OpenStudy (anonymous):

@TrojanPoem Can you help me with more?

OpenStudy (trojanpoem):

Of course.

OpenStudy (anonymous):

Solve the complex fraction x/x+3/1/x+1/x+3

OpenStudy (anonymous):

sorry that looks bad, I hate using the equation gadget on here

OpenStudy (anonymous):

okay Maggie. So we factorise quadratic expressions likes this. \[x ^{2}+3x-4\] You have to break the middle term in such a way that it adds up to get 3x again, but the broken parts multiply to give \[x ^{2} \times -4\]

OpenStudy (trojanpoem):

x/x+3/1/x+1/x+3 Looks horrible that way LOL

OpenStudy (trojanpoem):

Use the EQN and I will try to make you solve it not me.

OpenStudy (anonymous):

So it means I have to break my middle term to get \[-4x ^{2}\]. My middle term's 3x. I can break it into 4x - x This will add up to give 3x again, but will multiply to give -4x^2 Now you let the broken part be as it is and factorise by taking something common

OpenStudy (trojanpoem):

a sec

OpenStudy (trojanpoem):

If you want to simplify it just * (x)(x+3) and It will be (x^2) / (2x+3)

OpenStudy (anonymous):

Thanks! I may need more help, will you be online for a while?

OpenStudy (trojanpoem):

Not so long , like 10 min more.

OpenStudy (anonymous):

Okay, well thanks for your help!

OpenStudy (trojanpoem):

But did you understand how to solve it ?

OpenStudy (anonymous):

yes!